hhanan:

You have $1200 to invest in an account and need to have $1900 in one year. What interest rate would you need to have in order to reach this goal if the amount is compounded quarterly? Round your answer to the nearest percent.

hhanan:

options

CourtneyWatson:

Which one do you think is incorrect?

hhanan:

8

ScamTheMan:

Well, your first step is to figure out how you're going to make 1200 into 1900. Our first step would be to subtract 1900 by 1200 to figure out the difference. 1900 - 1200 = 700. Now that we have our number 700, we can use the basic formula for interest rate. Which is: I/PT=R I = Amount of Interest Generated P = Money Before Interest T = Time Period R = Interest Rate We substitute the numbers we have into our formula: 700/1200(4) = 0.145 We now take our answer and put it into percent form as per the question, making our answer: 14.5%. Since the only answer choice close to this is 14%, I'd just go ahead and choose that, there's always some margin of error.

hhanan:

thank you

ScamTheMan:

No problem, sorry I didn't get it before, I failed to see where it said "quarterly". I just had to adjust the formula and instead of putting "1", I used "4".